OpenStudy (anonymous):

if H(3)=1, H'(3)=3, F(3)=5, F'(3)=4, find (a) G'(3) if G(z)=F(z)*H(z)

OpenStudy (anonymous):

I assume you want to take the derivative of this, soo.... G(z)=F(z)*H(z) NOTE: []=PRIME [G(z)]=[F(z)](H(z))+(F(z))[H(z)] plug in the nubmers [G(3)]=(4)(1)+(5)(3) [G(3)]=19 I hope you can understand what I did there

OpenStudy (anonymous):

\[G'(3) = 5* 3 + 1 * 4 = 15+4 = 19\]

OpenStudy (anonymous):

Where did you find teh prime thingy?

OpenStudy (anonymous):

Remember to use the product rule. I showed the work, but it must have gotten deleted. The "prime thingy" is just one '

OpenStudy (anonymous):

Ohhh wait nevermind, I could have used '. LOLLL